# Wildfire Risk Prediction: A Survey of Recent Advances Using Deep Learning Techniques

### v4

@@ Abstract
Wildfires pose significant threats to ecosystems, wildlife, and human populations, leading to habitat destruction, pollutant emissions, and biodiversity loss. Accurate wildfire risk prediction is crucial for mitigating these impacts and protecting both environmental and human health. This paper presents a comprehensive review of wildfire risk prediction methodologies, particularly focusing on deep learning approaches. It begins by defining wildfire risk and summarizing the geographical distribution of relevant studies. Key predictive features, including meteorological, socio-economic, and fuel-related factors, are analyzed, with emphasis on feature collinearity evaluation and model interpretability to enhance predictive accuracy. Traditional methods such as wildfire danger rating systems and machine learning models are reviewed, followed by a categorization of deep learning models into time series prediction, image segmentation, and spatiotemporal prediction. Methods for translating model outputs into actionable probabilities are also discussed. The paper identifies the current challenges and limitations in existing models and outlines future opportunities for multimodal data integration and the development of architecturally efficient models for improved wildfire risk prediction.
@@ 1 Introduction
Wildfires occur in diverse environments such as forests, grasslands, farmlands, and peatlands, with significant impacts on global ecosystems, wildlife, human health, air quality, and carbon balance (Moritz et al., 2014; Stojanovic et al., 2016; Bowd et al., 2021; Holder et al., 2023; Lopez et al., 2024). They not only destroy plant communities, habitats, and human settlements but also contribute to elevated emissions of greenhouse gases and toxic pollutants like carbon dioxide, nitrogen oxides, methane, ozone, particulate matter, and mercury (Hua and Shao, 2017; Yang et al., 2022a; Pelletier et al., 2023; McNorton and Di Giuseppe, 2024). For instance, during the record-breaking 2023 wildfire season, approximately 18.5 million hectares burned across Canada (Erni et al., 2024), the largest area recorded since national reporting began in 1972. This season resulted in the evacuation of 232,000 people (Byrne et al., 2024), with over 200 communities evacuated, hundreds of homes, and infrastructure destroyed (Jain et al., 2024). The carbon emissions for this season were estimated at 647 Tg C (Byrne et al., 2024), with potential long-term impacts on communities and ecosystems expected to last for years or even decades. In the coming decades, climate change is expected to shorten fire return intervals and increase the burned area in vulnerable regions (Parks and Abatzoglou, 2020; Holdrege et al., 2024).
@@
In response to these challenges, wildfire risk prediction has attracted growing attention from governments, academia, and industry. Current prediction models consider essential components such as research scope delineation, data selection and processing, method establishment, setting prediction objectives, and evaluating model accuracy. However, a systematic review of these elements remains lacking. This paper addresses this gap by providing a comprehensive review of the data and deep learning methods used for wildfire risk assessment. In addition, the popular feature attribution techniques in deep-leaning-based wildfire risk prediction models are also reviewed. Finally, we also outline promising new deep learning methods.
@@
First, wildfire ignition and spread result from complex interactions among various factors, including fuel conditions, climate, weather, socio-economic factors, terrain, hydrology, and historical wildfire data (Fuller, 1991). This paper details the definitions, calculations, proxies, and frequency of use of these features, alongside an overview of publicly available spatial datasets for wildfire spread prediction.
@@
Second, A key concern in the design of prediction algorithms is evaluating feature collinearity and improving model interpretability. Including highly correlated features can degrade model performance and obscure feature importance (Li et al., 2022a), while model interpretability helps to clarify patterns of ignition and spread (Abdollahi and Pradhan, 2023). This paper introduces techniques for assessing feature collinearity and enhancing the interpretability of machine learning models.
@@
Another critical aspect of wildfire prediction is the algorithm establishment. Methods range from wildfire danger rating systems and statistical models to traditional machine learning and deep learning approaches. While Jain et al. (2020) provides a comprehensive review of machine learning models, this paper focuses on deep learning, which has seen rapid adoption and is becoming a crucial tool for wildfire predictions. To address the lack of systematic reviews in this area, we categorize deep learning-based models into three task types: time series prediction, image segmentation and classification, and spatiotemporal prediction. This study reviews existing deep learning models in these categories, providing a structured overview of their applications.
Finally, with the continuous advancements in deep learning techniques, the integration of multi-source geospatial data and deep learning models for wildfire risk prediction offers both significant opportunities and challenges. This paper also discusses the current limitations of existing algorithms and highlighting potential future developments in multimodal and architecturally efficient models.
@@ 2 Wildfire Risk Definition
In the majority of published wildfire risk methodology studies, the concept of wildfire risk is not clearly defined, leading to instances of confusion regarding the interchangeable use of terms such as risk, hazard, and danger. Consequently, this paper initiates by presenting a precise definition of wildfire risk. Miller and Ager (2012), based on the Society for Risk Analysis’s definition of risk, proposed a comprehensive fire risk framework (Fig. 1), which divides wildfire risk into three components: likelihood, intensity, and effects. This framework emphasizes the impact of wildfires on human activities and the environment rather than solely considering the probability of occurrence. For instance, low-intensity and low-impact wildfires might not significantly affect areas of concern, meaning that a high likelihood of wildfire occurrence does not necessarily equate to high wildfire risk.
@@
Additionally, their study distinguishes between commonly used terms in wildfire risk research, such as risk, hazard, exposure, threats, vulnerability, and fire danger. When summarizing related work, we focus only on studies concerning wildfire risk and danger, excluding those on hazard, exposure, and threats. The former primarily describes the outlook for fire occurrence over a certain period, while the latter tends to evaluate the hazard of wildfires (Fig.1). Consequently, if we consider only the likelihood of fire occurrence without considering its potential harm, the definition of wildfire risk can be simplified to the probability of a wildfire igniting or spreading at a specific location and time (Deeming, 1972). Many researchers in the past have described risk as the likelihood of wildfire occurrence, which also aligns with the wildfire framework established by the European Commission (Depicker et al., 2020).
@@
More specifically, these studies often treat wildfire risk assessment as a conditional probability, where wildfire risk is defined as the probability of a wildfire igniting or spreading under the influence of factors such as fuel, terrain, and weather conditions. In this context, wildfire risk equates to wildfire probability (Yang et al., 2004). In addition, some definitions emphasize the role of causal agents and their frequency in influencing wildfire risk (Hardy, 2005). In this definition, the importance of "causal agents" is highlighted—without these agents, the existing hazard would have zero potential to cause harm. Causal probability refers to the conditions that trigger wildfires, such as lightning or human activities Deeming (1972).
@@
Based on these definitions, the specific form of wildfire risk varies across studies. This paper categorizes wildfire risk definitions into four types based on the algorithms used and the research objectives: explicit definitions, implicit definitions, wildfire ignition risk, and wildfire burning risk. On one hand, definitions are categorized according to whether they explicitly represent various wildfire influencing factors and the contributions of historical wildfire records. These categories include definitions based on historical observed wildfire occurrence data, influencing factors, and simulation data, as well as data-driven machine learning implicit definitions. On the other hand, wildfire prediction research is generally divided into studies on wildfire ignition risk and wildfire burning risk. Wildfire ignition risk is typically estimated using ignition data, while wildfire burning risk refers to the risk of a fire encountering a particular place. Although this risk partly depends on ignition, it is also influenced by subsequent wildfire spread and fire suppression efforts and is generally derived from historical burn data or simulation methods (Miller and Ager, 2012).
@@ 2.1 Explicit and Implicit Definition Methods
Firstly, the simplest and most direct method in wildfire risk estimation is using historical fire occurrences to predict future risk. For example, Rubí and Gondim (2023) classified regions where one or more fires occurred within an hour as "high risk," while other areas were classified as "low risk." Preisler et al. (2004) selected all fire voxels within a past time ($t$) and spatial range and sampled a $\pi$ percentage of non-fire voxels using the scheme proposed by Brillinger et al. (2003). The occurrence of fire can then be characterized using conditional probability, which can be constrained within the (0, 1) interval using a sigmoid function.
$p(n)={\textstyle\sum_{t=0}^{T}}\frac{y_{t,n}}{T},\forall n\in[N]$ (1)
where $y_{t,n}$ is the wildfire occurrence at a specific time $t$ and location (grid cell) $n$ 0, 1; $N$ is the duration of the wildfire season. Similarly, Umunnakwe et al. (2022) estimated the spatiotemporal ignition probability using a similar function.
@@
where $m_{g,j}$ represents the fire ignition probability in grid cell $g$ during period $j$; $n_{g,j}$ and $N$ are the total number of wildfires and the total number of considered periods, respectively. They also introduced Monte Carlo techniques to interpolate grid cells. Lall and Mathibela (2016) assessed wildfire risk in Cape Town, South Africa, by defining daily risk ratings based on the average number of wildfires per day and their corresponding standard deviation over seven years. Specifically, fire occurrences below the average were rated as 1 or lower, those between the average and one standard deviation above were rated as 2 or moderate, those between one and two standard deviations above were rated as 3 or higher, and occurrences more than two standard deviations above the average were rated as 4 or extreme. An ANN model was then established using temperature, humidity, wind speed, and precipitation as inputs to predict wildfire risk.
The above models based on historical data do not consider the conditions for wildfire ignition or spread, thus having poor robustness and generalization ability. Moreover, due to the infrequent occurrence of wildfires over long time scales and large areas, there is a severe class imbalance between the occurrence and non-occurrence of wildfires in space and time. Therefore, the risk assessment results of these models are greatly influenced by data sampling. Consequently, some studies have begun to incorporate or solely use factors such as fuel, terrain, weather conditions, and wildfire triggers as proxies to estimate ignition risk. For example, Castro et al. (2003) used shrub moisture content, specifically live fuel moisture, as a proxy variable for wildfire risk in Spain. More complexly, Sadasivuni et al. (2013) defined wildfire risk as the interaction between human activity and fuel distribution. Specifically, they integrated population and fuel layers through additive operations as input variables for a gravity model in a spatial interaction model. The prediction results were then classified into five levels. Similarly, Gentilucci et al. (2024) categorized wildfire influencing factors and input them into a weight of evidence model, categorizing wildfire risk into low and high levels through regression of various influencing factor weights. Oliveira et al. (2021a); Bergonse et al. (2021) conducted classifications of land use, land cover, slope, elevation, and aspect. They then calculated the likelihood ratio for each category and determined fire susceptibility by summing the weight values of different categories. This process can be mathematically represented as follows:
@@
where $Lr_{i}$ represents the likelihood ratio of the variable corresponding to category $i$; $S_{i}$ and $S$ represent the number of burning pixels associated with category $i$ and the total number of burning pixels; $N_{i}$ and $N$ represent the number of pixels belonging to category $i$ and the total number of pixels in the study area; the total likelihood ratio $Lr_{j}$ of each grid cell $j$ is obtained by summing the $Lr_{i}$ values of all $n$ variables. Similarly, South Africa’s official wildfire danger rating system, the Lowveld model (Meikle and Heine, 1987; Lall and Mathibela, 2016), defines wildfire risk using the burn index (BI), wind correction factor (WCF), and rain correction factor (RCF), where BI is a function of temperature and relative humidity:
@@
where FDI represents the fire danger index. Srivastava et al. (2014) differed from the aforementioned studies by simultaneously considering both causal and anti-causal factors, including the distribution of check dams, firelines, and fire water towers, using an expert knowledge system to quantify the likelihood of forest fires.
@@
On the other hand, wildfire ignition and spread are highly random and exhibit a strong nonlinear relationship with various factors (Leuenberger et al., 2018). Data-driven machine learning methods are frequently employed in wildfire risk prediction research due to their capability to effectively model nonlinear functions using large-scale datasets (Jain et al., 2020). However, these methods do not explicitly quantify the contribution of individual factors to the overall wildfire risk, compared to explici definition methods. Thus, these methods can be categorized as implicit definition methods. Further details on these methods will be provided in Section 7.
@@
Apart from the impact of method selection on the prediction probability, these models also heavily depend on the construction of the label sample set. Different preprocessing techniques, such as historical resampling of wildfire data, can significantly affect the predicted probability values. Therefore, it is necessary to apply post-processing techniques, such as establishing a connection between explicit and implicit definitions, to enhance the interpretability and accuracy of the model outputs.
2.2 Wildfire Ignition and Burring Risk

### v5

# Deep Learning for Wildfire Risk Prediction: Integrating Remote Sensing and Environmental Data

@@ Abstract
Wildfires pose a significant threat to ecosystems, wildlife, and human communities, leading to habitat destruction, pollutant emissions, and biodiversity loss. Accurate wildfire risk prediction is crucial for mitigating these impacts and safeguarding both environmental and human health. This paper provides a comprehensive review of wildfire risk prediction methodologies, with a particular focus on deep learning approaches combined with remote sensing. We begin by defining wildfire risk and summarizing the geographical distribution of related studies. In terms of data, we analyze key predictive features, including fuel characteristics, meteorological and climatic conditions, socioeconomic factors, topography, and hydrology, while also reviewing publicly available wildfire prediction datasets derived from remote sensing. Additionally, we emphasize the importance of feature collinearity assessment and model interpretability to improve the understanding of prediction outcomes. Regarding methodology, we classify deep learning models into three primary categories: time-series forecasting, image segmentation, and spatiotemporal prediction, and further discuss methods for converting model outputs into risk classifications or probability-adjusted predictions. Finally, we identify the key challenges and limitations of current wildfire-risk prediction models and outline several research opportunities. These include integrating diverse remote sensing data, developing multimodal models, designing more computationally efficient architectures, and incorporating cross-disciplinary methods—such as coupling with numerical weather-prediction models—to enhance the accuracy and robustness of wildfire-risk assessments.
@@
[inst1]organization=Schulich School of Engineering, Department of Geomatics Engineering, University of Calgary,addressline=2500 University Dr NW, city=Calgary, postcode=T2N1N4, state=Alberta, country=Canada
@@
[inst2]organization=Department of Geography and Environmental Management, University of Waterloo,addressline=100 University Avenue, city=Waterloo, postcode=N2L3G1, state=Ontario, country=Canada
[…]
[inst3]organization=Department of Systems Design Engineering, University of Waterloo,addressline=100 University Avenue, city=Waterloo, postcode=N2L3G1, state=Ontario, country=Canada
[…]
[inst4] organization=CEREA, ENPC, EDF R&D, Institut Polytechnique de Paris, state=Ile-de-France, country=France
[…]
[inst5]organization=School of Emergency Management, Nanjing University of Information Science and Technology, addressline=219 Ningliu Road, city=Nanjing, postcode=210044, country=China
[…]
[inst6]organization=Division of Geoinformatics, KTH Royal Institute of Technology, addressline=Teknikringen 10a, cite=Stockholm, postcode=100 44, country=Sweden
[…]
[inst7]organization=School of Remote Sensing and Geomatics Engineering, Nanjing University of Information Science and Technology, addressline=219 Ningliu Road, city=Nanjing, postcode=210044, country=China
[…]
[inst8]organization=Department of Software Engineering, University of Waterloo,addressline=100 University Avenue, city=Waterloo, postcode=N2L3G1, state=Ontario, country=Canada
[…]
[inst9]organization=Department of Management Science and Engineering, University of Waterloo,addressline=100 University Avenue, city=Waterloo, postcode=N2L3G1, state=Ontario, country=Canada
@@ 1 Introduction
Wildfires are uncontrolled fires that occur in vegetated landscapes (Huidobro et al., 2024; Glossary of Wildland Fire Terminology, 2006). They can ignite naturally or be caused by human activities and are generally classified into planned fires, which are deliberately set for ecological or experimental purposes, and accidental fires, which result from natural causes or human negligence (Fillmore et al., 2024; Huidobro et al., 2024). Wildfires occur across diverse environments, including forests, grasslands, farmlands, and peatlands, where they play a significant role in shaping global ecosystems. They influence biodiversity, air quality, human health, and carbon balance, with both beneficial and detrimental effects (Moritz et al., 2014; Stojanovic et al., 2016; Bowd et al., 2021; Holder et al., 2023; Lopez et al., 2024). On the one hand, wildfires contribute to ecosystem dynamics by promoting nutrient cycling, facilitating species turnover, and enhancing vegetation regeneration (Bond et al., 2005; Pausas and Keeley, 2019). On the other hand, their uncontrolled occurrence poses substantial risks, including biodiversity loss, air pollution, threats to human settlements, and economic disruptions (Moritz et al., 2014; Bowd et al., 2021; Yang et al., 2022a; Lopez et al., 2024).
The magnitude of wildfire-related disasters has escalated in recent decades. For instance, the 2019–2020 "Black Summer" wildfires in Australia burned approximately 18.6 million hectares, caused 34 fatalities, destroyed over 5,900 structures, and resulted in an estimated economic loss exceeding \$110 billion (Naderpour et al., 2020; Haque et al., 2021). In North America, the 2018 Camp Fire in California led to 85 fatalities and the destruction of nearly 19,000 structures, while the 2023 Maui wildfires resulted in nearly 100 deaths and \$550 million in damages (Marris, 2023; Hertelendy et al., 2024). In Canada, wildfire frequency, size, and intensity have been increasing since the mid-20th century due to climate warming (Cunningham et al., 2024; Mulverhill et al., 2024). From 1959 to 2015, the total burned area and the number of large wildfires (over 200 hectares) have significantly risen, and the wildfire season has extended by approximately two weeks (Hanes et al., 2019; Jain et al., 2024). The 2023 wildfire season set a record, with 18.5 million hectares burned nationwide, displacing 232,000 people, destroying hundreds of homes, and releasing 647 Tg C into the atmosphere (Erni et al., 2024; Byrne et al., 2024; Jain et al., 2024). Investigations suggest that climate change will further shorten fire return intervals and expand burned areas in vulnerable regions (Parks and Abatzoglou, 2020; Holdrege et al., 2024).
@@
In response to these challenges, wildfire risk prediction has garnered increasing attention from governments, academia, and industry. Risk prediction involves delineating research scope, selecting and processing data, establishing methodologies, and understanding predictions. Despite the growing reliance on machine learning for wildfire risk prediction, a systematic review of these components remains lacking. This paper addresses this gap by providing a comprehensive review of the multi-source data and deep learning methods used in wildfire risk prediction, with a particular emphasis on feature attribution techniques and emerging deep learning models.
@@
First, wildfire ignition and spread result from complex interactions among multiple factors, including fuel conditions, climate, weather, socio-economic variables, terrain, hydrology, and historical wildfire occurrences (Fuller, 1991). This study systematically examines the definitions, proxies, frequency of use, and calculations (see Appendix) of these features, alongside an overview of publicly available datasets relevant to wildfire prediction.
@@
Second, an essential consideration in wildfire risk modeling is the evaluation of feature collinearity and model interpretability. Highly correlated features can obscure individual contributions and degrade model performance, especially for linear models (Li et al., 2022a), while model interpretability enhances the understanding of ignition and spread dynamics (Abdollahi and Pradhan, 2023). This paper introduces key techniques for assessing collinearity and improving interpretability in machine learning-based wildfire models.
@@
Third, the development of predictive algorithms varies widely, ranging from traditional wildfire danger rating systems and statistical models to modern machine learning and deep learning approaches. While previous studies have reviewed conventional machine learning techniques (Jain et al., 2020), this paper focuses on deep learning, which has seen rapid adoption and is emerging as a crucial tool for wildfire prediction. We categorize deep learning-based models into three primary task types: time-series forecasting, image segmentation and classification, and spatiotemporal prediction. A structured review of these models is presented, highlighting their applications and limitations.
Finally, integrating multi-source geospatial data with deep learning models for wildfire risk prediction presents both opportunities and challenges. This paper discusses the limitations of existing algorithms and explores potential advancements in multimodal and architecturally efficient models, as well as the possible potential of new data sources (such as microwave, LiDAR, and nightlight data) and cross-domain weather forecasting models. The fusion of these diverse datasets with deep learning techniques has the potential to enhance predictive accuracy and improve early warning systems. Given the increasing wildfire risk under climate change, developing robust, interpretable, and computationally efficient models is crucial for informing mitigation strategies, resource allocation, and adaptive management policies (Erni et al., 2024; Mulverhill et al., 2024).
@@ 2 Wildfire Risk Definition
In the majority of published wildfire risk methodology studies, the concept of wildfire risk is not clearly defined, leading to instances of confusion regarding the interchangeable use of terms such as risk, hazard, and danger. Consequently, this paper initiates by presenting a definition of wildfire risk. Miller and Ager (2012), based on the Society for Risk Analysis’s definition of risk, proposed a comprehensive fire risk framework (Fig. 1), which divides wildfire risk into three components: likelihood, intensity, and effects. This framework emphasizes the impact of wildfires on human activities and the environment rather than solely considering the probability of occurrence. For instance, low-intensity and low-impact wildfires might not significantly affect areas of concern, meaning that a high likelihood of wildfire occurrence does not necessarily equate to high wildfire risk. Notably, this framework does not explicitly incorporate the stochastic ignition caused by topographic and socioeconomic factors, which slightly diverges from the trends observed in the latest research.
@@
Additionally, their study distinguishes between commonly used terms in wildfire risk research, such as risk, hazard, exposure, threats, vulnerability, and fire danger. When summarizing related work, we focus only on studies concerning wildfire risk and danger, excluding those on hazard, exposure, and threats. The former primarily describes the outlook for fire occurrence over a certain period, while the latter tends to evaluate the hazard of wildfires (Fig 1).
Consequently, if wildfire risk is defined solely based on the likelihood of fire occurrence without accounting for its potential harm, it can be simplified to the probability of wildfire ignition or spread at a given location and time (Deeming, 1972). This interpretation aligns with the approach taken by many past studies, where risk has often been described as the probability of wildfire occurrence. Moreover, this definition is consistent with the wildfire risk framework established by the European Commission (Depicker et al., 2020).
@@
More specifically, these studies often treat wildfire risk assessment as a conditional probability, where wildfire risk is defined as the probability of a wildfire igniting or spreading under the influence of factors such as fuel, terrain, and weather conditions. In this context, wildfire risk equates to wildfire probability (Yang et al., 2004). In addition, some definitions emphasize the role of causal agents and their frequency in influencing wildfire risk (Hardy, 2005). The importance of "causal agents" is highlighted—without these agents, the existing hazard would have zero potential to cause harm. Causal probability refers to the conditions that trigger wildfires, such as lightning or human activities (Deeming, 1972).
@@
Based on these definitions, the specific form of wildfire risk varies across studies. This paper categorizes wildfire risk definitions into four types based on the algorithms used and the research objectives: explicit definitions, implicit definitions, wildfire ignition risk, and wildfire burning risk. On the one hand, definitions are categorized according to whether they explicitly represent various wildfire influencing factors and the contributions of historical wildfire records. These categories include explicit definitions based on historical observed wildfire occurrence data, influencing factors, and simulation data, as well as data-driven machine learning implicit definitions. On the other hand, wildfire prediction research is generally divided into studies on wildfire ignition risk and wildfire burning risk. Wildfire ignition risk is typically estimated using ignition data, while wildfire burning risk refers to the risk of a fire encountering a particular place. Although this risk partly depends on ignition, it is also influenced by subsequent wildfire spread and fire suppression efforts and is generally derived from historical burn data or simulation methods (Miller and Ager, 2012).
In this survey, it is important to note that the aforementioned definitions of wildfire risk differ from the concept of wildfire detection. Wildfire risk refers to the probability of a wildfire occurring, whereas wildfire detection focuses on identifying already ignited wildfires. Nevertheless, accurate wildfire detection can provide valuable insights for wildfire risk prediction, particularly by serving as historical fire occurrence data for constructing risk assessment or prediction models. Additionally, near-real-time wildfire detection (Hu et al., 2024; Zhang et al., 2024b) can offer crucial guidance for predicting fire spread and ignition risk.
@@ 2.1 Explicit and Implicit Definition Methods
Firstly, the simplest and most direct method in wildfire risk estimation is using historical fire occurrences to predict future risk. For example, Rubí and Gondim (2023) classified regions where one or more fires occurred within an hour as "high risk," while other areas were classified as "low risk." Similarly, Umunnakwe et al. (2022) estimated the spatiotemporal ignition probability using a similar function.
@@
where $m_{g,j}$ represents the fire ignition probability in grid cell $g$ during period $j$; $n_{g,j}$ and $N$ are the total number of wildfires and the total number of considered periods, respectively. Lall and Mathibela (2016) assessed wildfire risk in Cape Town, South Africa, by defining daily risk ratings based on the average number of wildfires per day and their corresponding standard deviation over seven years. Specifically, fire occurrences below the average were rated as 1 or lower, those between the average and one standard deviation above were rated as 2 or moderate, those between one and two standard deviations above were rated as 3 or higher, and occurrences more than two standard deviations above the average were rated as 4 or extreme. An artificial neural network (ANN) model was then established using temperature, humidity, wind speed, and precipitation as inputs to predict wildfire risk.
The above models based on historical data do not consider the conditions for wildfire ignition or spread, thus having poor robustness and generalization ability. Consequently, some studies have begun to incorporate or solely use factors such as fuel, terrain, weather conditions, and wildfire triggers as proxies to estimate ignition risk. Essentially, these methods can be defined as the contributions of various drivers to wildfire occurrence at specific times and locations. For example, Castro et al. (2003) used shrub moisture content, specifically live fuel moisture, as a proxy variable for wildfire risk in Spain. More complexly, the probability of wildfire occurrence can be defined as a conditional probability, representing the likelihood of a wildfire occurring given specific temporal, spatial, and environmental influencing factors. A typical example is the logistic model proposed by Preisler et al. (2004):
$p_{1}=\text{Prob}(N_{x,y,t}=1|U_{x,y,t})=\frac{\exp(\theta_{x,y,t})}{1+\exp(\theta_{x,y,t})}$ (2)
where $N_{x,y,t}$ and $U_{x,y,t}$ represent, respectively, whether a wildfire occurs and the collection of explanatory variables at location $(x,y)$ and time $t$. $\theta_{x,y,t}$ is a parameter estimated based on the explanatory variables, incorporating spatial and temporal dependencies. Considering the spatial, temporal, and nonlinear correlations of the explanatory variables, $\theta_{x,y,t}$ can be expressed as a combination of linear and nonlinear terms:
$\theta_{x,y,t}=\text{logit}(p_{1k})=$ $\ \text{fuel}_{k}+g_{1}(x_{k},y_{k})+g_{2}(\text{day}_{k})+g_{3}(\text{elev}_{k})$ (3)
$+g_{4}(\text{BI}_{k})+g_{5}(\text{FP}_{k})+g_{6}(\text{KB}_{k})+g_{7}(\text{TH}_{k})$
$+g_{8}(\text{Wspeed}_{k})+g_{9}(\text{RH}_{k})+g_{10}(\text{DBT}_{k})$
[…]
here, $g_{i}(\cdot)$ are nonparametric smoothing functions, typically estimated through generalized additive models, to capture nonlinear relationships. Variables such as $\text{fuel}_{k}$ and $\text{elev}_{k}$ represent explanatory factors, including fuel models, weather conditions (e.g., relative humidity, temperature), and fire danger indices (e.g., Burning Index (BI), Fire Potential Index (FP)), while $(x_{k},y_{k})$ are spatial coordinates, with $g_{1}(x_{k},y_{k})$ as a bivariate smoothing function for spatial nonlinear effects. In this context, $k$ denotes a unique spatiotemporal unit, corresponding to a specific location $(x,y)$ and time $t$, representing an individual voxel (e.g., a 1 km2 grid cell) and a specific day where wildfire occurrence probability is modeled.
Several other models have been employed to quantify the relationships between wildfire drivers and wildfire risk as well. For example, Sadasivuni et al. (2013) defined wildfire risk as the interaction between human activity and fuel distribution, integrating population and fuel layers through additive operations as inputs for a gravity model within a spatial interaction framework, with predictions classified into five risk levels. Similarly, Gentilucci et al. (2024) used a weight-of-evidence model to categorize wildfire drivers and classify wildfire risk into low and high levels based on regression of factor weights. Additionally, Oliveira et al. (2021a) and Bergonse et al. (2021) classified land use, land cover, slope, elevation, and aspect, calculated likelihood ratios for each category, and assessed fire susceptibility by summing weighted values across categories. This process can be mathematically expressed as follows:
@@
where $Lr_{i}$ represents the likelihood ratio of the variable corresponding to category $i$; $S_{i}$ and $S$ represent the number of burning pixels associated with category $i$ and the total number of burning pixels; $N_{i}$ and $N$ represent the number of pixels belonging to category $i$ and the total number of pixels in the study area; the total likelihood ratio $Lr_{j}$ of each grid cell $j$ is obtained by summing the $Lr_{i}$ values of all $n$ variables. Similarly, South Africa’s official wildfire danger rating system, the Lowveld model (Meikle and Heine, 1987; Lall and Mathibela, 2016), defines wildfire risk using the burning index, wind correction factor (WCF), and rain correction factor (RCF), where BI is a function of dry-bulb temperature (T) and relative humidity (RH):
@@
here, FDI refers to the fire danger index, while WCF and RCF are specified in Notice 1099 of 2013, issued in compliance with South Africa’s National Veld and Forest Fire Act (Department of Agriculture Forestry and Fisheries South Africa, 2013). Srivastava et al. (2014) differed from the aforementioned studies by simultaneously considering both causal and anti-causal factors, including the distribution of check dams, firelines, and fire water towers, using an expert knowledge system to quantify the likelihood of forest fires.
@@
On the other hand, wildfire ignition and spread are highly random and exhibit a strong nonlinear relationship with various factors (Leuenberger et al., 2018). Data-driven machine learning methods are frequently employed in wildfire risk prediction research due to their capability to effectively model nonlinear functions using large-scale datasets (Jain et al., 2020). While these methods are powerful for prediction, they do not explicitly quantify the contribution of individual factors to overall wildfire risk, unlike the explicit definition methods discussed earlier. As a result, they are classified as implicit definition methods. Further details on these methods will be provided in Section 6.
@@
In addition to the influence of method selection on predicted probabilities, these models are highly sensitive to the construction of the label sample set. Various preprocessing strategies—such as historical resampling of wildfire data—can substantially affect the resulting probability estimates. Therefore, it is essential to first establish a clear definition of wildfire risk rather than relying solely on binary labels. Furthermore, the application of post-processing techniques, such as linking explicit and implicit definitions of risk, can improve both the interpretability and accuracy of model outputs. This aspect will be further discussed in Section 6.1.
2.2 Wildfire Ignition and Burning Risk